OpenStudy (anonymous):

The slant height of the cone is 16 inches and the total area is 120 pi square inches, find the height of the cone?

OpenStudy (anonymous):

@amistre64

OpenStudy (anonymous):

r²+16r-120=0 r²+height²=16² using the above two equations h can be found out

OpenStudy (anonymous):

Basically 15 in.

OpenStudy (anonymous):

yeah i am with @sillynikki1 but exacally its The height is 14.83 inches.

OpenStudy (anonymous):

what is the formula of TA?

OpenStudy (anonymous):

@Hannah_Waller how did you get the r value?

OpenStudy (anonymous):

@Hannah_Waller here

OpenStudy (anonymous):

@sillynikki1

OpenStudy (anonymous):

how did you get h from this: r²+16r-120=0 r²+height²=16²

OpenStudy (anonymous):

SA = πr²+πrl 120π = πr²+16πr 0 = πr²+16πr - 120π r=5.56, r=-21.56 16² = h² + 5.56² (16²-5.56²)^(1/2) = h 15 = h